OpenStudy (keiraadianee):

The sum of what and 7x2 is 10x2.

OpenStudy (anonymous):

6? xD

OpenStudy (anonymous):

7*2 = 14 10*2 = 20 20-14= 6

OpenStudy (anonymous):

6 is your answer

OpenStudy (anonymous):